OpenStudy (anonymous):

An opinion poll asks an SRS of 1500 adults, "Do you happen to jog?" Suppose that the population proportion who jog (a parameter) is p = 0.15. To estimate p, we use the proportion p-hat in the sample who answer "Yes." Justify the use of a normal approximation and find the following probabilities

OpenStudy (anonymous):

a) P(p > 0.16) b) P(0.14 < p < 0.16)

OpenStudy (anonymous):

@Splash_dance can you help with this one also?

OpenStudy (anonymous):

First, why do you need to estimate the proportion of those who jog if you are given it in the beginning of your question? (i.e. you're told that p = 15%).

OpenStudy (anonymous):

For starters, we can use the Normal distribution (to approximate the Binomial distribution) because the sample size (n = 1500) is so large.

OpenStudy (anonymous):

would a be 0.1379

OpenStudy (anonymous):

So are you asking: Given that the population proportion, p, is equal to 0.15, what is the probability that p-hat (the result of your sample) is greater than 0.16?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, I might be way off but here it goes: First, the variance of a Bernoulli random variable is given by: \[\sigma ^{2} = p(1-p)\] In addition, when calculating distribution of the sample mean we use the formula: \[\frac{ sample mean - \mu }{ standard error of sample mean }\] So, to calculate the P(p>0.16), first we find the z-score associated with p=0.16: \[= \frac{ 0.16 - 0.15 }{ \sqrt{\frac{ 0.15\times0.85 }{ 1500 }} }\] which is approximately equal to 1.08465. Looking up 1.08 in a Z-score table we get: \[P(Z \le 1.08) = 0.8599\] Thus, the probability that Z is greater than 1.08 is: 1- 0.8599 = 14.01% Does this sound right to you?